The second of two numbers is seven more than the first. If the first in increased by 55, the result is five times the second. Find the numbers.
I tried this, but it is not working out: y=(x+7)=(x+55=5y)

Answer by fractalier(6550)

Call the numbers x and y.
From the facts in the problem
y = x + 7
and
x + 55 = 5y
Now substitute the first into the second and get
x + 55 = 5(x + 7)
x + 55 = 5x + 35
Collect like terms and get
20 = 4x
x = 5 and then
y = 12